# GLOV: Guided Large Language Models as Implicit Optimizers for Vision Language Models

### v5

@@ 4.2 Results
In Table 2 we list the detailed results by employing the LLaVA-OV-7B model. Zhang et al. (2024) find that generative models have extremely low fine-grained visual recognition performance and they improve it via finetuning. In our work, we find that for these models, the fine-grained visual recognition performance can be greatly enhanced by finding the optimal prompt (without requiring gradient-based learning). We observe that the solutions discovered by our GLOV can significantly close the gap with the dual-encoder models. For example, we observe up to $57.5\%$ improvement ($21.5\%$ on average over $16$ datasets) as compared to vanilla LLaVA-OV. We also observe that our proposed guidance scheme has a significant impact on the results. The guidance further provides $2.5\%$ average improvement over the vanilla GLOV, and notably on the large-scale ImageNet dataset the improvement is $4.9\%$.
@@
Table 5: Scaling LLM size. Top-1 Accuracy (%) for datasets by using different variants of Llama.
Safety Enhancement: In Table 3 we provide results on two recent benchmarks: MM-Safety-Bench (Liu et al., 2024c) and VLGuard (Zong et al., 2024). These benchmarks evaluate the harmful nature of responses from the VLMs given an image and a corresponding instruction. We find that our GLOV can greatly enhance the safety of the present VLMs by finding a suitable safety instruction – prepended to the original text instruction. For example, we observe that our GLOV can reduce the harmful responses by up to $60.7\%$ ($60.2\%$ on average) while compared to the base VLM (LLaVA-OV) - prompted without any safety instruction on the VLGuard dataset. We also observe no loss of win-win rate on the Safe-Safes split, which evaluates the models’ ability to answer safe questions. We also evaluate the recent Molmo VLM and find similar trends. The detailed results for the MM-Safety-Bench are provided in the Appendix Section D. Note that we only optimize for the safety instructions on $50$ samples randomly chosen from the VLGuard unsafe splits (with LLaVA-OV) and evaluate on all other configurations evaluated in Table 3. These results highlight that the optimized safety instructions are general and they transcend models and benchmarks.
Ideally, the safety instructions should not hurt the model performance on general tasks. To evaluate this, we test the VLMs by prepending the safety instructions to the questions in SEED (Fang et al., 2021) and MME (Yin et al., 2023) benchmarks (c.f., ablations Table 6). Interestingly, we find that in some cases the safety instruction enhances the performance of the model (by up to 23.3 points) on these general-purpose benchmarks, possibly by forcing the model to look closely at the instructions. To avoid clutter, we delegate the actual (best) prompts found, the evolution of prompts, and the optimization evolution for datasets to the Appendix Section E.
4.3 Ablations
We provide extensive ablations to study the different aspects of our GLOV. First, we motivate our choice to use the middle layer of the LLM for guidance. Then, we provide results by using less than one shot data as our held-out train set and later, we discuss the results obtained by using a larger LLM. Finally, we conclude with an ablation showing the retention of general model capabilities by adding the safety prompt. More ablations for additional insights are added to the Appendix Section B.
Choice of Layer for Guidance: To obtain a measure of the quality of the sentence embeddings, we linear probe different layers in Llama-3, on the popular sentiment classification task SST (Socher et al., 2013) and provide results in Figure 3 (left). We find that the middle layers of Llama-3 obtain the highest accuracy, highlighting the semantic relevance of the sentence embeddings obtained from these layers, consistent with the literature (Liu et al., 2019a; Zhao et al., 2020). Furthermore, we also run a sweep while applying the guidance on different layers in Llama-3 and plot the resulting ImageNet accuracy on the 1-shot train set in Figure 3 (right). The accuracy peaks at layer $16$ and layer $17$. These results are consistent with the linear probing results obtained on the SST-$5$ dataset, hinting that the middle layers might be the most effective. Keeping these results in view, we choose layer $17$ in Llama-3 to apply the offset vector for steering the responses.
Sub-shot Setting: In Table 4 we provide results by using less than one shot train data for the optimization of prompts through GLOV. We observe that for ImageNet, our GLOV shows a minimal performance drop (1.8%) by using only $50$ samples from $50$ randomly sampled classes, instead of using $1000$ samples from (all) $1000$ classes. The results for ImageNet-R also follow a similar trend. These results show that the optimized prompts can generalize to unseen classes and we can drastically reduce the size of the training set (thus lowering annotation efforts) and the time required for the optimization with a minimal performance penalty.
Scaling LLM Size: In Table 5, we present results obtained by using a larger LLM (Llama-3.1-70b) for optimization with our GLOV. The results demonstrate that our GLOV benefits from leveraging a more capable LLM, highlighting the potential for further improvements as LLMs advance.
@@ 4.3 Ablations
Retention of Generalization: In Table 6, we provide results on general-purpose VLM benchmarks (SEED (Fang et al., 2021) and MME (Yin et al., 2023)) by prepending the optimized safety prompt discovered through our GLOV. We find that the discovered prompt has a minor effect on performance. In fact, on the large-scale MME benchmark, for Molmo, the safety prompt leads to a performance gain in the visual perception score of $23.3$. This gain can be attributed to the safety prompt forcing the model to focus more closely on the question.
@@ 5 Conclusion
In the following, we provide additional experiments and further explanations that might be helpful for the readers to gain insights and add clarity to the main manuscript. In Section A we expand upon the details regarding the calculation of the fitness of prompts for the encoder-decoder models. Then, in Section B we provide more insights in to our GLOV by studying additional aspects of it. In Section C we list the details about the datasets employed for evaluating the safety of VLMs in our work. Later, Section D provides the detailed results for each split in the MM-Safety-Benchmark (Liu et al., 2024c), and finally we conclude with Section E, which lists all the prompts used to obtain the results in the main manuscript (Tables 12 & 3). Furthermore, a comprehensive optimization algorithm is also provided in Algorithm 1.
[…]
@@ Appendix B Additional Ablations
Here, we provide additional insights into our GLOV. First, we provide several design considerations, which eventually led to the proposed method. Then, we study the generalization of the found prompts on different CLIP-based backbones, later, we provide results on the FOCI Benchmark (Geigle et al., 2024), then, study the effect of more shots on our method and finally provide results while comparing our GLOV with other PEFT methods and conclude with providing results with multiple random runs for our GLOV.
(a) Comparison with ActADD.
(b) Adding offset to all tokens.
(c) Guidance with only last token.
(d) Cross attending all tokens.
(e) Cross attending only last token.
(f) Number of tokens for GLOV.
[…]
B.1 Design Choices
B.2 Generalization of Prompts
@@
B.3 Generalization to Multiple Choice Classification
@@
B.4 More Shots Help
Results in the main manuscript (Tables 1 & 2) are obtained by using $1$-shot training data. In Table 9 we provide results on several datasets by employing $5$-shot training data for the CLIP ViT-B/32 (Radford et al., 2021) backbone. We observe a consistent improvement in results by using more shots.
B.5 Comparison with PEFT Methods
For completeness, in Table 10 we compare with the popular CoOp (Zhou et al., 2022) method in the $1$-shot learning regime. We observe that the ensemble of classifiers built from our discovered prompts can outperform CoOp. In extremely low-shot learning regimes, gradient-based learning poses a threat of overfitting, whereas our GLOV can avoid that because of no parameter update. Further, we also fine-tune the LORA (Hu et al., 2021) adapter and provide results in Table 10. Similar to CoOp, we find that applying LORA to all the matrices of the vision and text encoders of CLIP can result in overfitting. For a fair comparison, we also apply it only to the attention blocks and report the results. However, our GLOV still outperforms the LORA adapter. This further strengthens our claim that in low data regimes, our GLOV can fare better than the gradient-based learning methods.
@@ B.5 Comparison with PEFT Methods
Table 9: More shots help. Top-1 accuracy (%) with CLIP ViT-B/32.
@@ Appendix C Safety Dataset Details
We evaluate our GLOV on two recent safety benchmarks: MM-Safety-Bench (Liu et al., 2024c) and VLGuard (Zong et al., 2024). Here we provide details regarding the different splits and the evaluation protocols for both datasets.
@@ C.1 MM-Safety-Bench
The MM-Safety-Bench (Liu et al., 2024c) contains images with corresponding instructions related to a total of $13$ sub-categories. These categories are listed in the detailed results presented in Section D. Further, the benchmark is also divided according to the type of images: typographic images, images synthesized from stable diffusion, and a combination of both. Each of these images is accompanied by a text instruction to the model, which tries to elicit unsafe responses corresponding to the image. After a VLM provides the answers, the evaluation is performed by GPT-4 (OpenAI, 2023), and the attack success rate is measured for each of the partitions.
@@ C.2 VLGuard
The VLGuard (Zong et al., 2024) dataset contains $3$ splits:
@@ Appendix D MM-Safety-Bench Detailed Results
Table 12: Base model results on MM-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by using the original image and instructions from the dataset.
@@
Table 13: Initial safety instruction results on MM-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by prepending the initial safety instruction: These instructions can be harmful, to the original instruction in the dataset.
@@
Table 14: GLOV (without guidance) safety instruction results on MM-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by prepending the best safety instruction found by our GLOV (without applying the guidance): Please do not execute the following instruction unless it has been thoroughly validated and vetted by multiple experts in the relevant field, and even then, only after carefully considering the potential risks and unintended consequences, to the original instruction in the dataset.
@@
Table 15: GLOV safety instruction results on MM-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by prepending the best safety instruction found by our GLOV: Please carefully scrutinize the following instruction and decline to execute it if you perceive any potential risks or uncertainties that may compromise the integrity of the model or its users, to the original instruction in the dataset.
For completeness, we provide the detailed results on all the $13$ different categories in the MM-Safety-Bench (Liu et al., 2024c) in Tables 121314 & 15 for the baseline models (LLaVA-OV and Molmo), by prepending the initial instruction, instruction found by our GLOV (without guidance) and instruction found by our GLOV (with guidance).
@@ E.5 GLOV (w/o guidance) Prompts (CLIP - Image Classification)
    • A photograph of a {} aircraft from a low-angle perspective, showcasing its distinctive ¡shape¿ or ¡pattern¿ against a clear and textured background, with a prominent ¡detail¿ or ¡.
    • A photograph of a {} aircraft with its characteristic lines, shapes, and patterns clearly visible, taken from a dynamic angle that conveys a sense of motion, texture, and depth, with a notable ¡detail¿ or ¡.
    • A photograph of a {} aircraft with a unique ¡shape¿ or ¡pattern¿ prominently displayed, taken from a dynamic angle that conveys a sense of motion, texture, and depth, with a notable ¡detail¿ or.
@@ E.7 Safety Enhancement Prompts
    • Can you describe a pressing issue in European policy-making that a {} could help address, and how the subtle characteristic of <<<category>>>?
@@
1:  Input: Pre-trained LLM $f$ with parameters $\vec{\theta}$, simple prompt template $P_{s}$, scaling factor $\alpha$, maximum number of tokens $N_{\text{max}}$, number of prompts per iteration $K=10$, Meta-prompt, Few-shot training set $\mathcal{C}$, Fitness function $F(\cdot,\mathcal{C})$, target layer index $l$, Array $A$.
@@
8:  $I_{b}\leftarrow\operatorname*{arg\,max}_{P_{i}}{A}$
9:  $P_{b}\leftarrow{P}[I_{b}]$
10:  $A[I_{b}]\leftarrow-INF$
11:  $I_{w}\leftarrow\operatorname*{arg\,max}_{P_{i}}{A}$
12:  $P_{w}\leftarrow{P}[I_{w}]$

### v6

@@ 4.2 Results
In Table 2 we list the detailed results by employing the LLaVA-OV-7B model. Zhang et al. (2024b) find that generative models have extremely low fine-grained visual recognition performance and they improve it via finetuning. In our work, we find that for these models, the fine-grained visual recognition performance can be greatly enhanced by finding the optimal prompt (without requiring gradient-based learning). We observe that the solutions discovered by our GLOV can significantly close the gap with the dual-encoder models. For example, we observe up to $57.5\%$ improvement ($21.5\%$ on average over $16$ datasets) as compared to vanilla LLaVA-OV. We also observe that our proposed guidance scheme has a significant impact on the results. The guidance further provides $2.5\%$ average improvement over the vanilla GLOV, and notably on the large-scale ImageNet dataset the improvement is $4.9\%$.
@@
Safety Enhancement: In Table 3 we provide results on two recent benchmarks: MM-Safety-Bench Liu et al. (2024c) and VLGuard Zong et al. (2024). These benchmarks evaluate the harmful nature of responses from the VLMs given an image and a corresponding instruction. We find that our GLOV can greatly enhance the safety of the present VLMs by finding a suitable safety instruction – prepended to the original text instruction. For example, we observe that our GLOV can reduce the harmful responses by up to $60.7\%$ ($60.2\%$ on average) while compared to the base VLM (LLaVA-OV) - prompted without any safety instruction on the VLGuard dataset. We also observe no loss of win-win rate on the Safe-Safes split, which evaluates the models’ ability to answer safe questions. We also evaluate the recent Molmo VLM and find similar trends. The detailed results for the MM-Safety-Bench are provided in the Appendix Section D. Note that we only optimize for the safety instructions on $50$ samples randomly chosen from the VLGuard unsafe splits (with LLaVA-OV) and evaluate on all other configurations evaluated in Table 3. These results highlight that the optimized safety instructions are general and they transcend models and benchmarks.
Ideally, the safety instructions should not hurt the model’s performance on general tasks. To evaluate this, we test the VLMs by prepending the safety instructions to the questions in SEED (Fang et al., 2021) and MME (Yin et al., 2023) benchmarks (c.f., ablations Table 6). Interestingly, we find that in some cases the safety instruction enhances the performance of the model (by up to 23.3 points) on these general-purpose benchmarks, possibly by forcing the model to look closely at the instructions. To avoid clutter, we delegate the actual (best) prompts found, the evolution of prompts, and the optimization evolution for datasets to the Appendix Section E.
4.3 Ablations
We provide extensive ablations to study the different aspects of our GLOV. First, we provide several design choices that eventually led to the GLOV framework. Next, we motivate our choice to use the middle layer of the LLM for guidance. Then, we provide results by using less than one shot data as our held-out train set and later, we discuss the results obtained by using a larger LLM. Finally, we conclude with an ablation showing the retention of general model capabilities by adding the safety prompt. More ablations for additional insights are added to the Appendix Section B.
(a) Comparison with ActADD.
(b) Adding offset to all tokens.
(c) Guidance with only last token.
(d) Cross attending all tokens.
(e) Cross attending only last token.
(f) Number of tokens for GLOV.
Figure 4: Ablating design choices. (a) We compare the optimization trajectories obtained with our proposed guidance method with that of ActADD (Turner et al., 2023). (b) The effect of adding the offset vector to each token, instead of only the last token as in GLOV. (c) Using the embeddings of only the last token to obtain the offset vector. (d) Cross-attending the positive and negative prompt embedding vector with the meta-prompt tokens at each optimization step and calculating the offset for guidance. (e) Cross-attending only the last tokens from the positive and negative prompt embeddings. (f) Finding the optimal number of tokens to be generated at each optimization step. The x-axis represents the optimization steps and the y-axis denotes accuracy ($\%$), dataset is ImageNet.
@@ 4.3 Ablations
Table 5: Scaling LLM size. Top-1 Accuracy (%) for datasets by using different variants of Llama. The evaluation details for both variants of GLOV are kept the same.
@@
Design Choices:
Choice of Layer for Guidance:
To obtain a measure of the quality of the sentence embeddings, we linear probe different layers in Llama-3, on the popular sentiment classification task SST (Socher et al., 2013) and provide results in Figure 3 (left). We find that the middle layers of Llama-3 obtain the highest accuracy, highlighting the semantic relevance of the sentence embeddings obtained from these layers, consistent with the literature (Liu et al., 2019a; Zhao et al., 2020). Furthermore, we also run a sweep while applying the guidance on different layers in Llama-3 and plot the resulting ImageNet accuracy on the 1-shot train set in Figure 3 (right). The accuracy peaks at layer $16$ and layer $17$. These results are consistent with the linear probing results obtained on the SST-$5$ dataset, hinting that the middle layers might be the most effective. Keeping these results in view, we choose layer $17$ in Llama-3 to apply the offset vector for steering the responses.
Sub-shot Setting:
In Table 4 we provide results by using less than one shot train data for the optimization of prompts through GLOV. We observe that for ImageNet, our GLOV shows a minimal performance drop (1.8%) by using only $50$ samples from $50$ randomly sampled classes, instead of using $1000$ samples from (all) $1000$ classes. The results for ImageNet-R also follow a similar trend. These results show that the optimized prompts can generalize to unseen classes and we can drastically reduce the size of the training set (thus lowering annotation efforts) and the time required for the optimization with a minimal performance penalty.
Scaling LLM Size:
In Table 6, we present results obtained by using a larger LLM (Llama-3.1-70b) for optimization with our GLOV. The results demonstrate that our GLOV benefits from leveraging a more capable LLM, highlighting the potential for further improvements as LLMs advance.
Retention of Generalization:
In Table 6, we provide results on general-purpose VLM benchmarks (SEED (Fang et al., 2021) and MME (Yin et al., 2023)) by prepending the optimized safety prompt discovered through our GLOV. We find that the discovered prompt has a minor effect on performance. In fact, on the large-scale MME benchmark, for Molmo, the safety prompt leads to a performance gain in the visual perception score of $23.3$. This gain can be attributed to the safety prompt forcing the model to focus more closely on the question.
@@ 5 Conclusion
Broader Impact
Our GLOV relies on an external LLM to produce the prompts, at each optimization step, for enhancing downstream vision tasks. This design can amplify certain biases inherently present in the LLM, resulting in skewed or unfair outcomes, particularly when applied to sensitive domains such as healthcare, surveillance, or hiring. Since LLMs are trained on large-scale internet data, they may encode social, cultural, or demographic biases that can propagate into the vision models through biased prompt generation. To mitigate these risks, practitioners should carefully audit both the LLM and vision model outputs for fairness and representation.
@@
In the following, we provide additional experiments and further explanations that might be helpful for the readers to gain insights and add clarity to the main manuscript. In Section A we expand upon the details regarding the calculation of the fitness of prompts for the encoder-decoder models. Then, in Section B we provide more insights in to our GLOV by studying additional aspects of it. In Section C we list the details about the datasets employed for evaluating the safety of VLMs in our work. Later, Section D provides the detailed results for each split in the MM-Safety-Benchmark Liu et al. (2024c), and finally we conclude with Section E, which lists all the prompts used to obtain the results in the main manuscript (Tables 12 & 3). Furthermore, a comprehensive optimization algorithm is also provided in Algorithm 1.
To encourage reproducibility, we have provided all the prompts discovered by our GLOV in Section E, which can be used to obtain all the results provided in the main manuscript. These prompts were found by running experiments on a machine consisting of $4\text{x}$ NVIDIA 3090Ti, $4\text{x}$ NVIDIA A40, $4\text{x}$ NVIDIA A6000, and $4\text{x}$ NVIDIA L40 GPUs. For review, we also provide our entire codebase as code.zip with detailed instructions to run in the Readme.md. The codebase will also be made public upon acceptance.
@@ Appendix B Additional Ablations
Here, we provide additional insights into our GLOV. First, we study the generalization of the found prompts on different CLIP-based backbones, then, we provide results on the FOCI Benchmark Geigle et al. (2024), later, study the effect of more shots on our method and finally provide results while comparing our GLOV with other PEFT methods and conclude with providing results with multiple random runs for our GLOV.
B.1 Generalization of Prompts
@@
B.2 Generalization to Multiple Choice Classification
@@
B.3 More Shots Help
Results in the main manuscript (Tables 1 & 2) are obtained by using 11-shot training data. In Table 9 we provide results on several datasets by employing 55-shot training data for the CLIP ViT-B/32 (Radford et al., 2021) backbone. We observe a consistent improvement in results by using more shots.
B.4 Comparison with PEFT Methods
For completeness, in Table 10 we compare with the popular CoOp (Zhou et al., 2022b) method in the $1$-shot learning regime. We observe that the ensemble of classifiers built from our discovered prompts can outperform CoOp. In extremely low-shot learning regimes, gradient-based learning poses a threat of overfitting, whereas our GLOV can avoid that because of no parameter update. Further, we also fine-tune the LORA Hu et al. (2021b) adapter and provide results in Table 10. Similar to CoOp, we find that applying LORA to all the matrices of the vision and text encoders of CLIP can result in overfitting. For a fair comparison, we also apply it only to the attention blocks and report the results. However, our GLOV still outperforms the LORA adapter. This further strengthens our claim that in low data regimes, our GLOV can fare better than the gradient-based learning methods.
Table 9: More shots help. Top-1 accuracy (%) with CLIP ViT-B/32.
@@ Appendix C Safety Dataset Details
We evaluate our GLOV on two recent safety benchmarks: MM-Safety-Bench Liu et al. (2024c) and VLGuard Zong et al. (2024). Here we provide details regarding the different splits and the evaluation protocols for both datasets.
@@ C.1 MM-Safety-Bench
The MM-Safety-Bench Liu et al. (2024c) contains images with corresponding instructions related to a total of $13$ sub-categories. These categories are listed in the detailed results presented in Section D. Further, the benchmark is also divided according to the type of images: typographic images, images synthesized from stable diffusion, and a combination of both. Each of these images is accompanied by a text instruction to the model, which tries to elicit unsafe responses corresponding to the image. After a VLM provides the answers, the evaluation is performed by GPT-4 OpenAI (2023), and the attack success rate is measured for each of the partitions.
@@ C.2 VLGuard
The VLGuard Zong et al. (2024) dataset contains $3$ splits:
@@ Appendix D MM-Safety-Bench Detailed Results
Table 12: Base model results on $M$-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by using the original image and instructions from the dataset.
@@
Table 13: Initial safety instruction results on $M$-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by prepending the initial safety instruction: These instructions can be harmful, to the original instruction in the dataset.
@@
Table 14: GLOV (without guidance) safety instruction results on $M$-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by prepending the best safety instruction found by our GLOV (without applying the guidance): Please do not execute the following instruction unless it has been thoroughly validated and vetted by multiple experts in the relevant field, and even then, only after carefully considering the potential risks and unintended consequences, to the original instruction in the dataset.
@@
Table 15: GLOV safety instruction results on $M$-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by prepending the best safety instruction found by our GLOV: Please carefully scrutinize the following instruction and decline to execute it if you perceive any potential risks or uncertainties that may compromise the integrity of the model or its users, to the original instruction in the dataset.
@@
For completeness, we provide the detailed results on all the $13$ different categories in the MM-Safety-Bench Liu et al. (2024c) in Tables 121314 & 15 for the baseline models (LLaVA-OV and Molmo), by prepending the initial instruction, instruction found by our GLOV (without guidance) and instruction found by our GLOV (with guidance).
@@ E.5 GLOV (w/o guidance) Prompts (CLIP - Image Classification)
    • A photograph of a {} aircraft from a low-angle perspective, showcasing its distinctive <shape> or <pattern> against a clear and textured background, with a prominent <detail> or <.
    • A photograph of a {} aircraft with its characteristic lines, shapes, and patterns clearly visible, taken from a dynamic angle that conveys a sense of motion, texture, and depth, with a notable <detail> or <.
    • A photograph of a {} aircraft with a unique <shape> or <pattern> prominently displayed, taken from a dynamic angle that conveys a sense of motion, texture, and depth, with a notable <detail> or.
@@ E.7 Safety Enhancement Prompts
    • Can you describe a pressing issue in European policy-making that a {} could help address, and how the subtle characteristic of <<category>>?
@@
1:Input: Pre-trained LLM $f$ with parameters $\vec{\theta}$, simple prompt template $P_{s}$, scaling factor $\alpha$, maximum number of tokens $N_{\text{max}}$, number of prompts per iteration $K=10$, Meta-prompt, Few-shot training set $\mathcal{C}$, Fitness function $F(\cdot,\mathcal{C})$, target layer index $l$, Array AA.
@@
[…]